this is my question please solve and send it in to my mail id nitin_sharma38@rediffmail.com please
Q"- A aeroplane flies 3500 km at the same time train travel 600 km. if the aeroplane flies 50 km/h more than 5 times of the speed of the train than find the speed of the aeroplane and the train?

Answer by mananth(16946)

A aeroplane flies 3500 km at the same time train travel 600 km. if the aeroplane flies 50 km/h more than 5 times of the speed of the train than find the speed of the aeroplane and the train?
let the speed of the train be x km/hr
Time taken by train = 600/x hours
Speed of plane = 5x+50
Time taken by plane = 3500/ 5x+50
The times for both are same
So
600/x= 3500 / 5x+50
3500x=600(5x+50)
3500x=3000x+30000
500x= 30000
x=60
.
60kms/hr is the speed of the train.
Ananth